# A Survey on Vision-Language-Action Models for Embodied AI

### v4

@@ I Introduction
Figure 1: The general architecture of vision-language-action models. Important related components are shown in the dashed-line boxes.
@@
  • Resources. We summarize the necessary resources for training and evaluating VLA models, including recent datasets and benchmarks in real-world or simulated environments. We also discuss various approaches to address current challenges, such as data scarcity and inconsistency.
@@ II Background
II-A Unimodal Models
Vision-language-action models integrate three modalities, often relying on existing unimodal models. The transition from convolutional neural networks (e.g., ResNet [16]) to visual Transformers (e.g., ViT [17], SAM [18]) in computer vision has facilitated the development of vision foundation models (VFMs). In natural language processing, the evolution from recurrent neural networks (e.g., LSTM [19], GRU [20]) to Transformers [4] initially led to the pretrain-finetune paradigm (e.g., BERT [21], ChatGPT [1]), followed by the recent success of prompt tuning driven by large language models. Reinforcement learning (e.g., DQN [5], AlphaGo [22], PPO [23], Dactyl [24]) has also witnessed a shift towards employing Transformers (§III-A1) to model the Markov Decision Process as autoregressive sequential data.
II-B Vision-Language Models
Vision-language tasks, encompassing image captioning [25], visual question answering [26], visual grounding [27], require the fusion of computer vision and natural language processing models. Early efforts, such as Show and Tell [28], leveraged the success of early CNNs and RNNs. The advent of high-capacity language models, including BERT [21] and GPT [29], ushered in a new era of Transformer-based VLMs. One of the pioneering self-supervised pretraining methods is ViLBERT [30], while CLIP [31] popularized contrastive pretraining approaches for multimodal alignment. The recent emergence of large language models has driven the development of multimodal LLMs (MLLMs) or large multimodal models (LMMs), which achieve state-of-the-art performance on multimodal instruction-following tasks. Representative MLLMs include Flamingo [32], BLIP-2 [7], and LLaVA [8]. VLMs share a close relationship with VLAs, as the multimodal architectures of VLMs can be readily adopted for VLAs. Additionally, VLMs can function as high-level task planners if they possess sufficient reasoning capabilities.
II-C Embodied AI & Robot Learning
@@ III-A1 Reinforcement Learning
Reinforcement learning (RL) trajectories, characterized by state-action-reward sequences, naturally align with the structure of sequence modeling problems, rendering them well-suited for Transformer models. These RL Transformers have become foundational to many recent VLA models. Pioneering efforts in this direction include Decision Transformer (DT) [34] and Trajectory Transformer (TT) [35]. Gato [36] further extends this paradigm to a multi-modal/task/embodiment setting.
Reinforcement learning from human feedback (RLHF) has become an important training ingredient for LLMs. SEED [37] utilizes RLHF alongside skill-based reinforcement learning to address the sparse reward issue of long-horizon tasks in robot learning. Reflexion [38] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
@@ III-A2 Pretrained Visual Representations
Table I: Pretrained visual representations. V: Vision. L: Language. CL: contrastive learning. TFM: Transformer. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation. For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc. $\mathcal{S}(\cdot)$ is some similarity measurement.
Model V-Encoder Type Objective Notation Robotic Tasks
CLIP [31] ViT-B [17] VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair (Used by CLIPort [43], EmbCLIP [44], and CoW [45])
R3M [39] ResNet-50 [16] Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Sim-Mani: Meta-World, Franka Kitchen, Adroit
MVP [41] ViT-B
ViT-L
MAE [46] $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch Real-Mani (xArm7): pick, reach block, push cube, close fridge
@@
VC-1 [47] ViT-L MAE, CL (A study of previous works, CLIP, R3M, MVP, VIP.) Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0; Sim-Navi: Habitat
Voltron [48] TFM [4] MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
@@
RPT [42] ViT [17, 41] MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x,y,z$: three distinct modalities Real-Mani (Franka): stack, pick, pick from bin
@@
I-JEPA [52] ViT JEPA $\frac{1}{M}\sum_{i=1}^{M}\sum_{j}\left\lVert x_{j}(i)-y_{j}(i)\right\rVert_{2}^{2}$ $x_{j}(i)$: $j$-th masked patch embedding of block $i$; $y_{j}(i)$: unmasked patch embedding -
Theia [53] ViT-T/S/B Distillation (Distillation of vision foundation models, including ViT, CLIP, SAM, DINOv2, and Depth-Anything.) Sim-Mani&Navi: CortexBench [47]; Real-Mani: pick, place, open door/drawer
@@ III-A4 Dynamics Learning
Model V-Encoder Type Objective Notation Robotic Tasks
Vi-PRoM [61] ResNet Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$: predicts frame index in raw seq. $x$ given shuffled seq. $x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open and close drawer/door
MaskDP [62] ViT-MAE [46] MAE (implicit fwd & inv dynamics) $\sum_{i=1}^{T}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: state or action token; $y$: the other modality Sim: DeepMind Control Suite
MIDAS [63] ViT, Mask R-CNN [64] Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: VIMA-Bench
@@
$+\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)\big{)}$
@@
PACT [66] ResNet-18, PointNet [67] Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Navi: Habitat; Real-Navi (MuSHR vehicle)
@@
GR-1 [69] ViT-MAE Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: CALVIN
@@ III-A8 Reasoning
ThinkBot [90] applies it to recover missing action descriptions in sparse human instructions. RAT [91] combines retrieval-augmented generation and CoT to elicit context-aware reasoning, improving long-horizon generation. ReAct [92] interleaves verbal reasoning traces and actions to allow for diverse types of reasoning. Tree-Planner [80] employs a tree-of-thoughts approach. ECoT [93] re-trains OpenVLA [50] using reasoning chain datasets, enabling embodied CoT reasoning that further enhances its performance.
Pros and Cons.
@@
Table III: Low-level control policies. We also include some non-VLA models as they are closely related, marked in gray text. BC: behavior cloning (action type, cont/disc: continuous/discrete). TFM: Transformer. Xattn: cross-attention. Concat: concatenation. $p$/$s$: prompt/state vision encoder. [SC]: self-collect data. MPC: model predictive control.
Model V-Encoder L-Encoder VL-Fusion Action-Decoder Objective Training Dataset Robotic Tasks
CLIPort [43] CLIP-ResNet50, Transporter-ResNet CLIP-GPT Hadamard LingUNet BC ($\textbf{SE}(2)$) [SC] Sim: Ravens [94]
MCIL [95] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP, [SC] Sim: 3D Playroom environment
HULC [96],
HULC++ [97]
MCIL CNN Sentence-BERT LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
Language costs [98] CLIP-ViT,
UNet [99]
CLIP-GPT Concat STORM Max Likelihood Est. (cost map) [SC] Sim & Real (Franka): pick, place
Interactive Language [100] ResNet CLIP-GPT Xattn TFM BC (cont) Language-Table [SC] Sim & Real (xArm6): rearrangement
Hiveformer [101] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data Sim: RLBench [102]
PerAct [103] 3D CNN CLIP-GPT Concat PerceiverIO 3D affordance RLBench data Sim: RLBench; Real (Franka): put marker, place food, etc.
Act3D [104] CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
RVT, RVT-2 [105, 106] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data, [SC] Sim: RLBench; Real (Franka): stack, press, place
RoboPoint [107] ViT-L/14 Vicuna-V1.5 13B [108] Concat - 2D affordance (project to 3D) [SC] Real (Franka): pick, place
Gato [36] ViT Sent.Piece Concat TFM BC (cont & disc) [SC] Sim & Real (Sawyer): RGB-stacking
RoboCat [109] VQ-GAN ($p,s$) [110] - - TFM BC, Observation prediction Self-improve Sim & Real (Sawyer, Franka, KUKA): stacking, building, lifting, insertion, removal
VIMA [111] ViT, Mask R-CNN [64] T5 [112] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data [SC] Sim (Ravens): VIMA-Bench
BC-Z [113] ResNet18 ($p,s$) USE FiLM MLP BC (cont) [SC] Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
RT-1 [114] EfficientNet [115] USE FiLM TFM BC (disc) RT-1-Kitchen [SC] Real (Everyday Robots): pick-place, move, knock
MOO [111] OWL-ViT ($p$) [116], EfficientNet ($s$) USE FiLM TFM BC (disc) [SC] Real (Everyday Robots): pick, move near, knock, place upright, place into
Q-Transformer [117] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1-Kitchen Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
RT-Trajectory [118] EfficientNet - - TFM BC (disc) [SC] Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
ACT [119] ResNet18 - - CVAE-TFM BC (cont, action chunking) [SC] with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
MT-ACT [120] (RoboAgent) CNN - FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
RoboFlamingo [121] CLIP-ViT-L/14 LLaMA [122], MPT [123] Xattn LSTM, TFM BC (cont) CALVIN data Sim: CALVIN
VoxPoser [125] ViLD [126],
MDETR [127],
OWL-ViT [116],
SAM [18]
GPT-4 Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
UniPi [128] Imagen video [129] T5-XXL Xattn CNN Inverse dynamics PF [SC], Ravens, BridgeV1 Sim: Painting Factory (PF), Ravens
Diffusion Policy [130] ResNet18 - - UNet,
TFM
DDPM [131] [SC] Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
DP3 [132] DP3 Encoder (Model design follows Diffusion Policy) DDIM [SC] Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
SUDD [133] ResNet18 CLIP-GPT Concat UNet,
TFM
DDPM Lang-Guided Data Gen Sim: MuJoCo; Real (UR5e): pick, place
Octo [134] CNN T5-base Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
MDT [136] CLIP-ViT-B/16 ($p$), Voltron/ResNet18 ($s$) CLIP-GPT Concat DiT [137] DDIM [138] CALVIN, LIBERO, [SC] Sim: CALVIN, LIBERO; Real (Franka): pick, push, open, close, etc.
RDT-1B [139] SigLIP T5-XXL Xattn DiT DDPM (Aggregated Datasets) Real (ALOHA dual-arm): wash, pour, fold, etc.
RT-2 [2] ViT-22B [140],
ViT-4B [141]
PaLI-X, PaLM-E Concat Symbol-tuning BC (disc), Co-fine-tuning VQA, RT-1-Kitchen Real: RT-1 evaluation tasks, Language-Table
RT-H [142] ViT-22B [140] PaLI-X 55B Concat Symbol-tuning BC (disc) Diverse+Kitchen Real: Diverse+Kitchen eval tasks
RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE [SC] Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
OpenVLA [50] DINOv2, SigLIP [144] Prismatic-7B [145] Concat Symbol-tuning BC (disc) OXE, DROID [146] Real: BridgeV2, RT-1 evaluation tasks
$\pi_{0}$ [147] SigLIP PaliGemma Concat Action expert Flow matching OXE, $\pi$-cross-embod. [SC] Real (general robot control): sweep, open, pack, etc.
@@ III-B Low-level Control Policies
The diversity among control policy networks arises from individual modules and overall architectures. The most typical architecture is shown in Figure 1. This section explores various approaches to designing low-level control policies. Table III summarizes their technical details.
@@ III-B1 Non-Transformer Control Policies
CLIPort [43] integrates the vision and language encoders of CLIP [31] with the Transporter network [94], creating a two-stream architecture. In one stream, CLIP’s vision encoder extracts “semantic” information from the RGB image, while in the other stream, the Transporter network extracts “spatial” information from the RGB-D image. The CLIP sentence encoder encodes the language instruction and guides the output action, which is a pair of end-effector poses: the picking and placing poses. CLIPort demonstrates the ability to pick and place objects as specified by the language instruction.
BC-Z [113] processes two types of task instructions: a language instruction or a human demonstration video. The environment is presented to the model in the form of an RGB image. Then the instruction embedding and the image embedding are combined through the FiLM layer [148], culminating in the generation of actions. This conditional policy is asserted to exhibit zero-shot task generalization to unseen tasks.

### v5

@@ I Introduction
Figure 1: The general architecture of vision-language-action models. Three representative methods for action prediction are shown. Related components are presented in dashed boxes.
@@
  • Resources. We summarize the necessary resources for training and evaluating VLAs, including recent datasets and benchmarks in real-world or simulated environments. We also discuss various approaches to address current challenges, such as data scarcity and inconsistency.
@@ III-A1 Reinforcement Learning
Reinforcement learning (RL) laid the foundation for embodied AI and continues to help advance VLAs. Deep Q-Network (DQN) first demonstrated learning policies directly from high-dimensional pixel inputs, underscoring the need for greater model capacity in end-to-end RL. RL trajectories—sequences of states, actions, and rewards—naturally align with sequence modeling problems, making them well-suited to transformer architectures. Pioneering efforts in this direction include Decision Transformer (DT) [17] and Trajectory Transformer (TT) [18]. Gato [19] further extended this paradigm to a multimodal, multitask, multi-embodiment setting. These methods in RL directly inspired modern VLAs [20, 21], which also process multimodal sequences using transformer-based architectures.
Furthermore, a synergy has emerged between RL and LLMs, benefiting embodied AI in multiple ways. Reinforcement learning from human feedback (RLHF) aligns LLMs with human preferences and has also been applied to robot learning. SEED [22] utilizes RLHF alongside skill-based RL to address the sparse reward issue and to improve safety via human evaluation. Conversely, LLMs also enable novel RL methods: Reflexion [23] proposes a novel verbal reinforcement learning framework that replaces weight updates in RL models with linguistic feedback, which is naturally applicable to embodied decision making where planning also occurs in language form. Eureka [24] shows that LLMs can design reward functions for embodied agents that outperform expert human-engineered ones. As LLMs evolve, this synergy accelerates progress in RL and embodied AI.
@@ III-A2 Pretrained Visual Representations
Table I: Pretrained visual representations. V: Vision. L: Language. CL: contrastive learning. TFM: Transformer. Sim/Real: simulated/real-world. Mani/Navi: manipulation/navigation. For simplicity, we only show the main part of the objective equation, omitting elements such as temperature, auxiliary loss, etc. $\mathcal{S}(\cdot)$ denotes a similarity measure.
Model Network Type Objective Notations Robotic Tasks
CLIP [25] ViT-B VL-CL $\sum_{i=1}^{N}-\log\frac{\exp\left(\mathcal{S}(x_{i},y_{i})\right)}{\sum_{j=1}^{N}\exp\left(\mathcal{S}(x_{i},y_{j})\right)}$ $(x_{i},y_{i})$: image-text pair (Used by CLIPort [31], EmbCLIP [32], and CoW [33])
R3M [26] ResNet-50 Time-CL $\sum_{i=1}^{T}-\log\frac{\exp\left(\mathcal{S}(x_{i},x_{j})\right)}{\exp\left(\mathcal{S}(x_{i},x_{j})\right)+\exp\left(\mathcal{S}(x_{i},x_{k})\right)}$ $x_{i}$: $i$-th video frame, $i<j<k$ Sim-Mani: Meta-World, Franka Kitchen, Adroit
MVP [28] ViT-B/L MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i}\right)$ $x_{i}$: image patch Real-Mani (xArm7): pick, reach block, push cube, close fridge
@@
VC-1 [34] ViT-L MAE, CL (A combination of previous works, including CLIP, R3M, MVP, and VIP.)   Sim-Mani: Meta-World, Adroit, DMC, TriFinger, Habitat 2.0; Sim-Navi: Habitat
Voltron [35] TFM MAE, Lang-Gen $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$;
@@
RPT [30] ViT MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x,y,z$: three distinct modalities Real-Mani (Franka): stack, pick, pick from bin
@@
I-JEPA [39] ViT JEPA $\frac{1}{M}\sum_{i=1}^{M}\sum_{j}\left\lVert x_{j}(i)-y_{j}(i)\right\rVert_{2}^{2}$ $x_{j}(i)$: $j$-th masked patch embedding of block $i$; $y_{j}(i)$: unmasked patch embedding
Theia [40] ViT-T/S/B Distillation (Distillation of vision foundation models: ViT, CLIP, SAM, DINOv2, Depth-Anything.)   Sim-Mani&Navi: CortexBench (VC-1); Real-Mani: pick, place, open door/drawer
@@ III-A4 Dynamics Learning
Model Vision Encoder Type Objective Notations Robotic Tasks
Vi-PRoM [48] ResNet Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$: predicts frame index in raw seq. $x$ given shuffled seq. $x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open & close drawer/door
MaskDP [49] ViT (MAE) MAE (implicit fwd & inv dynamics) $\sum_{i=1}^{T}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: state or action token; $y$: the other modality Sim: DeepMind Control Suite
MIDAS [50] ViT, Mask R-CNN Inverse dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: VIMA-Bench
@@
    $\operatorname{MSE}\left(a_{t},f_{\text{inv}}(s_{t},s_{t+1})\right)\big{)}$
@@
PACT [52] ResNet-18, PointNet Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Navi: Habitat; Real-Navi (MuSHR vehicle)
@@
GR-1 [54] ViT (MAE) Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: CALVIN
@@ III-A8 Reasoning
Reasoning is naturally compatible with high-level task planning, as both often occur in the language domain. ThinkBot [75] applies CoT to recover missing action descriptions in sparse human instructions, thereby enhancing instruction coherence and boosting success rates in difficult tasks. ReAct [76] interleaves reasoning traces and actions, where CoT can help create action plans, inject commonsense knowledge, and handle exceptions, ultimately improving embodied decision-making. RAT [77] integrates CoT with retrieval-augmented generation (RAG) to mitigate hallucinations and thus improve embodied planning. Tree-Planner [65] employs a tree-of-thoughts approach for task planning.
Another paradigm involves equipping low-level control policies with reasoning capabilities, as pioneered by ECoT [78]. This method trains OpenVLA [37] to conduct embodied CoT reasoning about plans, sub-tasks, motions, and visual features, before predicting low-level actions. By relying on this multi-step reasoning rather than the “muscle memory” of VLAs, it improves success rates on challenging generalization tasks without requiring additional robot data.
Strengths and Limitations.
@@ III-B Low-level Control Policies
It can also be referred to as a low-level policy, low-level controller, or action primitive. The diversity among VLAs arises from individual modules and overall architectures. The general architecture is shown in Figure 1. This section explores various approaches to designing low-level control policies. Table III summarizes their technical details.
@@ III-B1 Non-Transformer Control Policies
CLIPort [31] integrates CLIP with the Transporter network, creating a two-stream architecture. The CLIP vision encoder extracts “semantic” information from the RGB image, while the Transporter network extracts “spatial” information from the RGB-D image. The CLIP sentence encoder encodes the language instruction and guides the output $SE(2)$ action, consisting of paired pick and place end-effector poses. It represents an early demonstration of language-conditioned pick-and-place capabilities.
BC-Z [79] processes two types of task instructions: a language instruction or a human demonstration video. The environment is presented to the model in the form of an RGB image. Then the instruction embedding and the image embedding are combined through the FiLM layer, culminating in the generation of actions. This conditional policy is asserted to exhibit zero-shot task generalization to unseen tasks.
@@
Table III: VLAs that serve as low-level control policies. $\lozenge$ indicates large VLAs. Closely related (non-VLA models) are included in brackets. The remaining are generalized VLAs. BC: behavior cloning (cont/disc: continuous/discrete action). Xattn: cross-attention. Concat: concatenation. Quant.: quantization. pp/$s$: prompt/state vision encoder. [SC]: self-collect data.
Model Vision
Encoder
Language Encoder / VLM Action Decoder/Head Architecture Training Objectives Training Datasets Environments, Embodiments, Tasks, and Skills
CLIPort [31] CLIP-ResNet50, Transporter-ResNet CLIP-GPT LingUNet Hadamard BC ($SE(2)$) [SC] Sim: Ravens
MCIL [80] Custom-CNN USE RNN LMP MCIL Play-LMP, [SC] Sim: 3D Playroom environment
HULC [81, 82] MCIL CNN Sentence-BERT RNN LMP MCIL CALVIN data Sim: CALVIN
Language costs [84] CLIP-ViT, UNet CLIP-GPT STORM Concat MLE (cost map) [SC] Sim & Real (Franka): pick, place
Interactive Language [85] ResNet CLIP-GPT TFM Xattn BC (cont) [SC: Language-Table] Sim & Real (xArm6): rearrangement
Hiveformer [86] UNet CLIP-GPT TFM, UNet Concat BC (cont) RLBench data Sim: RLBench
PerAct [87] 3D CNN CLIP-GPT PerceiverIO Concat 3D affordance RLBench data, [SC] Sim: RLBench; Real (Franka): pick, place, stack, open drawer, sweep, insert peg, etc.
Act3D [88] CLIP-ResNet50 CLIP-GPT TFM Xattn 3D affordance (voxel) RLBench data, [SC] Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
RVT [89],
RVT-2 [90]
CLIP-ResNet50 CLIP-GPT TFM Concat 2D affordance (project to 3D) RLBench data, [SC] Sim: RLBench; Real (Franka): stack, press, place
RoboPoint [91] ViT-L/14 Vicuna-V1.5 13B   Concat 2D affordance (project to 3D) [SC] Real (Franka): pick, place
Gato [19] ViT Sent.Piece TFM Concat BC (cont & disc) [SC] Sim & Real (Sawyer): RGB-stacking
(RoboCat) [92] VQ-GAN ($p,s$)   TFM Quant. BC, observation prediction Self-improvement Sim & Real (Sawyer, Franka, KUKA): stacking, building, lifting, insertion, removal
VIMA [93] ViT, Mask R-CNN T5 TFM Xattn BC ($SE(2)$) [SC:VIMA-Data] Sim (Ravens): VIMA-Bench
BC-Z [79] ResNet18 ($p,s$) USE MLP FiLM BC (cont) [SC] Real (EDR): pick-place/wipe/drag, grasp, push
RT-1 [94] EfficientNet USE TFM FiLM BC (disc) [SC: Fractal] Real (EDR): pick-place, move, knock
MOO [93] OWL-ViT ($p$), EfficientNet ($s$) USE TFM FiLM BC (disc) [SC] Real (EDR): pick, move near, knock, place upright, place into
Q-Transformer [95] EfficientNet USE TFM FiLM TD error Fractal, Auto-collect Sim: pick; Real (EDR): pick, place, open/close drawer, move near
(RT-Trajectory) [96] EfficientNet   TFM   BC (disc) [SC] Real (EDR): pick, place, fold towel, swivel chair, etc.
(ACT) [97] ResNet18   CVAE-TFM   BC (cont, action chunking) [SC] with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
RoboAgent / MT-ACT [98] CNN   CVAE-TFM FiLM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
RoboFlamingo [99] CLIP-ViT-L/14 LLaMA, MPT, GPT-NeoX LSTM, TFM Xattn BC (cont) CALVIN data Sim: CALVIN
DeeR-VLA [101] (Model design follows RoboFlamingo++) BC (cont) CALVIN data Sim: CALVIN
Instruct2Act [102] CLIP, SAM ChatGPT Robot APIs Tool-Use     Sim (Ravens): VIMA-Bench
VoxPoser [103] ViLD, MDETR, OWL-ViT, SAM GPT-4 MPC Tool-Use     Sim: Sapien; Real (Franka): move&avoid, set table, close drawer, open bottle, sweep trash
UniPi [83] Imagen video T5-XXL CNN Xattn Inverse dynamics [SC:PF], Ravens, BridgeV1 Sim: Painting Factory (PF), Ravens
(Diffusion Policy) [104] ResNet18   UNet, TFM   DDPM [SC] Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
(DP3) [105] DP3 Encoder (Model design follows Diffusion Policy) DDIM [SC] Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
SUDD [106] ResNet18 CLIP-GPT UNet, TFM Concat DDPM Lang-guided data generation Sim: MuJoCo; Real (UR5e): pick, place
Octo [107] CNN T5-base TFM Concat DDPM OXE Real: BridgeV2, CMU Baking, Stanford Coffee, Berkeley Peg Insert, etc.
MDT [109] CLIP-ViT-B/16 ($p$), Voltron/ResNet18 ($s$) CLIP-GPT DiT Concat DDIM CALVIN, LIBERO, [SC] Sim: CALVIN, LIBERO; Real (Franka): pick, push, open, close, etc.
RDT-1B [110] SigLIP T5-XXL DiT Xattn DDPM (Aggregated Datasets) Real (ALOHA dual-arm): wash, pour, fold, etc.
RT-2 [2] $\lozenge$ ViT-4B, ViT-22B PaLI-X, PaLM-E Symbol-tuning Concat BC (disc), Co-fine-tuning Fractal, VQA Sim: Language-Table; Real: RT-1 evaluation tasks
RT-H [111] $\lozenge$ (Model design follows RT-2) BC (disc) Diverse+Kitchen Real: Diverse+Kitchen eval tasks
RT-X [112] $\lozenge$ (Models from RT-1 and RT-2) BC (disc) [SC: OXE] Real: BridgeV2, RT-1 evaluation tasks, etc.
OpenVLA [37] $\lozenge$ DINOv2, SigLIP Prismatic-7B Symbol-tuning Concat BC (disc) OXE, DROID Real: BridgeV2, RT-1 evaluation tasks, Franka-Tabletop, DROID, etc.
OpenVLA-OFT [113] $\lozenge$ (Improves OpenVLA with OFT recipe) BC (cont, parallel decode w/ chunk.) LIBERO, [SC] Sim: LIBERO; Real (ALOHA setup): fold, scoop, put
TraceVLA [114] $\lozenge$ (Model design follows OpenVLA, adding visual trace prompting) BC (disc) BridgeV2, Fractal, [SC] Sim: SimplerEnv; Real (WidowX): pick, push, fold, swipe
$\pi_{0}$ [115] $\lozenge$ SigLIP PaliGemma Action expert Concat Flow matching OXE, [SC: $\pi$-cross-embod.] Real (general robot control): sweep, open, pack, etc.
RoboMamba [116] $\lozenge$ CLIP-ViT Mamba MLP Concat BC (cont) [SC] Sim: Sapien; Real (Franka): open door/box, staple
SpatialVLA [117] $\lozenge$ SigLIP, Ego3D Position Encoding PaliGemma 2 MLP Concat BC (Adaptive Action Grids) OXE, BridgeV2, Fractal, RH20T Sim: SimplerEnv, LIBERO; Real (WidowX): pick, place, close, push
TinyVLA [118] $\lozenge$ ViT Pythia Diffusion Policy Concat DDPM OXE, [SC] Sim: MetaWorld; Real (Franka, UR5): place, stack, flip mug, close drawer, open box
CogACT [119] $\lozenge$ DINOv2, SigLIP LLaMA 2 DiT Concat DDIM OXE, [SC] Sim: SimplerEnv; Real (Realman, Franka): pick, place, stack, open/close oven
DexVLA [20] $\lozenge$ ViT, ResNet-50 Qwen2-VL 2B, DistilBERT, ScaleDP Concat DDPM [SC] Real (Franka, UR5e, AgileX): pick, fold shirt, bus table, pour
HybridVLA [120] $\lozenge$ DINOv2, SigLIP, CLIP LLaMA 2, Phi-2 MLP Concat DDIM OXE, DROID Sim: RLBench; Real (Franka, AgileX): pick-place, pour, open drawer, fold, etc.
LAPA [121] $\lozenge$ C-ViViT, VQ-GAN LWM-Chat-1M MLP Concat LAPA OXE, BridgeV2, Something V2 Sim: Language-Table, SimplerEnv; Real (Franka): pick, cover with towel, knock over
WorldVLA [122] $\lozenge$ VQ-GAN Chameleon TFM Quant. BC (disc, chunk.), world model LIBERO, OpenVLA data Sim: LIBERO
UniVLA [123] $\lozenge$ MoVQ Emu3 TFM Quant. BC (disc, chunk.), world model (Sim benchmark data) Sim: CALVIN, LIBERO, SimplerEnv